OpenStudy (anonymous):

A machine is set up to cut metal strips of varying lengths and widths based on the time (t) in minutes. The change in length is given by the function l(t)=t2−t√, and the change in width is given by w(t)=t2−2t12. Which function gives the change in area of the metal strips?

OpenStudy (anonymous):

oh your new to this

OpenStudy (anonymous):

If you could help me that would be great (:

OpenStudy (anonymous):

yea sure i can try

OpenStudy (anonymous):

is it multiple choice

OpenStudy (anonymous):

yes a(t)=t4+2t a(t)=t4+2t+3t52 a(t)=t4−3t52+2t a(t)=t4+2t−2t12+t√

OpenStudy (anonymous):

so what is your idea first

OpenStudy (anonymous):

if you have one